The magnitude of earthquakes in a certain region in North America can be modeled as an exponential distribution with mean 2.7, as measured on the Richter scale. Find the probability that exactly two of the next 5 earthquakes will exceed 6.0 on the Richter scale (assume the earthquakes’ magnitudes are independent from each other).
Answer #1

X ~ exp (λ = 1 / 2.7 = 0.3704 )
P ( X > x ) = 1 - ( 1 - e(-λ* x) )

X ~ B ( n = 5 , P = 0.1084 )